OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \text{prove that if a and b are positive integers } \hspace{.33em}\\~\\ \text{then } \sqrt 2\hspace{.33em}\\~\\ \text{lies between} \hspace{.33em}\\~\\ \dfrac{a}{b}~~~\text{and}~~~\dfrac{a+2b}{a+b} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (zarkon):

do two cases... case 1) \(\dfrac{a}{b}>\sqrt{2}\) case 2) \(\dfrac{a}{b}<\sqrt{2}\) and use the fact that \[\frac{a+2b}{a+b}=\frac{a+b+b}{a+b}=\frac{a+b}{a+b}+\frac{b}{a+b}=1+\frac{b}{a+b}\]

OpenStudy (mathmath333):

still confused on how to use this fact and proceed

OpenStudy (michele_laino):

please note that if a/b<sqrt(2), the we have: \[\frac{ a }{ b }+1<\sqrt{2}+1\] and: \[\frac{ 1 }{ \frac{ a }{ b }+1 }>\frac{ 1 }{ \sqrt{2}+1 }=\sqrt{2}-1\] so: \[1+\frac{ b }{ a+b }=1+\frac{ 1 }{ \frac{ a }{ b }+1 }>1+\sqrt{2}-1\] or: \[1+\frac{ b }{ a+b }=1+\frac{ 1 }{ \frac{ a }{ b }+1 }>\sqrt{2}\]

OpenStudy (zarkon):

Assume \(\dfrac{a}{b}>\sqrt{2}\) \[\frac{a+2b}{a+b}=1+\frac{b}{a+b}=1+\frac{1}{\frac{a}{b}+1}<1+\frac{1}{\sqrt{2}+1}=\sqrt{2}\]

ganeshie8 (ganeshie8):

Here is a complete proof using same argument as above Let \(p = \frac{a}{b}\) and \(q = \dfrac{p+2}{p+1}\), then we need to show \(\sqrt{2}\) lies between \(p\) and \(q\). Let \(A\) be the set of all positive rationals \(p\) such that \(p^2 \lt 2\) and let \(B\) be the set of all positive rationals \(p\) such that \(p^2\gt 2\). If \(p\) is in \(A\) then \(p^2 \lt 2 \). \(q^2-2 = \left(\dfrac{p+2}{p+1}\right)^2 - 2 = \dfrac{2-p^2}{(p+1)^2} \gt 0\tag{1}\) If \(p\) is in \(B\) then \(p^2 \gt 2 \). \(q^2-2 = \left(\dfrac{p+2}{p+1}\right)^2 - 2 = \dfrac{2-p^2}{(p+1)^2} \lt 0\tag{2}\) the result follows from above two inequalities. \(\blacksquare\)

OpenStudy (mathmath333):

thanks all !!